Please help me solve this equation: If you divide a number by 6, then subtract 5, then add 3, you get 5

Answer by ikleyn(52781)

.
Please help me solve this equation: If you divide a number by 6, then subtract 5, then add 3, you get 5
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Make all these operations in the reverse order:

Distract 3 from 5:     (5-3 = 2)

then add 5:            (2+5 = 7)

then multiply by 6:    (6*7 = 42)

Answer.  42.